# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

### v1

@@ 8.2. Potential relationships to open problems for PDEs
Let $\{\nu_{\varepsilon,\delta}\}_{\varepsilon,\delta>0}$ be a family of ergodic invariant measures of (1.3). Inspired by [BC17] and [CP22], the large deviations for $\{\nu_{\varepsilon,\delta}\}_{\varepsilon,\delta>0}$ under a scaling regime $(\varepsilon,\delta(\varepsilon))\rightarrow(0,0)$ strongly rely on uniform dynamical large deviations, and its rate function is expected to be governed by the quasi-potential. Based on the discussion in Remark 8.4, one can only expect to observe the rate function for the lower bound governed by the restricted quasi-potential (8.4). This connects to the open problem of uniqueness for the SQG equation (8.5).
As long as one can prove large deviations for $\{\nu_{\varepsilon,\delta(\varepsilon)}\}_{\varepsilon>0}$ with rate function (8.4), then Proposition 8.3 indicates the following relationships. On the one hand, if a solution of (8.5) exists in $\mathcal{A}_{0}$ (which implies uniqueness), then, by applying Proposition 8.3, one can observe that the expected rate function (8.4) equals to the Gaussian rate function. Conversely, if non-Gaussian large deviations can be proven for $\{\nu_{\varepsilon,\delta(\varepsilon)}\}_{\varepsilon>0}$ in the sense of Definition 8.5, the uniqueness problem of (8.5) could be answered negatively. Therefore, Proposition 8.3 suggests a new perspective for understanding the uniqueness problem of the PDE (8.5) through probabilistic approaches. However, rigorously establishing this relationship requires proving large deviations in equilibrium, which lies beyond the scope of this paper.
@@
It is worth pointing out that in [Tot20], the existence of Gaussian stationary solutions for (1.3) with $\delta=0$ is proved using the energy solutions approach. However, this does not contradict the discussion of non-Gaussian large deviations for invariant measures in the sense of Definition 8.5, as the aforementioned non-Gaussian large deviations focus on fluctuations in spaces with higher regularity due to the regularization of the noise and the choice of the scaling regime $(\varepsilon,\delta(\varepsilon))$.
On the other hand, in [HLZZ24], it was shown that there exist infinitely many non-Gaussian ergodic stationary solutions of (1.3) with $\delta=0$, which lie in the Besov space $B^{-1/2}_{p,1}$ for any $p\geqslant 2.$ In their work, the parameters $\alpha$ and $\beta$ satisfy $\alpha\in[0,3/4)$ and $\beta\in[0,1/8).$ We hope to bridge the discussion between non-Gaussian ergodic stationary solutions and the uniqueness problem of the PDE (8.5) via the non-Gaussian large deviations.
@@ Appendix A A Examples of the Weak-strong Uniqueness Classes
In this section, we provide some examples of the weak-strong uniqueness classes $\mathcal{C}_{0}$ and their $\mathcal{I}$-closure $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$ for different choices of $\alpha$ and $\beta$ satisfying (1.2). These examples cover subcritical, critical and supercritical cases.
@@
Suppose that $11/2<\alpha<$ 1 and $\beta=\alpha/2+1/4$. Let
$\mathcal{C}_{0}^{1}=\{\theta\in\mathbb{X}_{\alpha,\alpha/2+1/4}:\theta\in L^{r}\left([0,T];L^{p}(\mathbb{T}^{2})\right),\,\frac{1}{p}+\frac{\alpha}{r}=\alpha-\frac{1}{2},\,p\geqslant 1,\,r>0\}.$
@@
$\frac{1}{2}\|\theta_{2}(t)\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\|\theta_{2}(s)\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\theta_{2}(0)\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\langle\Lambda^{\alpha-1}\theta_{2},g\rangle\,\mathrm{d}s.$ (A.1)
@@
The same conclusion holds in the case where $\beta=\alpha/2$, which was stated in [CW99, Theorem 2.2, Remark 2.3]. For every $t\in[0,T]$, we expand $\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}$ as
$\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}=2\langle\Lambda^{-1/2}\theta_{1}(t),\Lambda^{-1/2}\theta_{2}(t)\rangle+\|\theta_{1}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}+\|\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}$ (A.2)
$\quad\int_{0}^{t}\left(2\langle\Lambda^{\alpha-1/2}\theta_{1},\Lambda^{\alpha-1/2}\theta_{2}\rangle+\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}\theta_{1}\rangle+\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}\theta_{2}\rangle\right)\,\mathrm{d}s$ (A.3)
$=-\langle\Lambda^{-1/2}\theta_{2}(t),\Lambda^{-1/2}\theta_{1}(t)\rangle+\langle\Lambda^{-1/2}\theta_{2}(0),\Lambda^{-1/2}\theta_{1}(0)\rangle+\int_{0}^{t}\langle\Lambda^{\alpha-1/2}(\theta_{2}+\theta_{1}),g\rangle\,\mathrm{d}s.$
For any $\epsilon>0$, let $\eta_{\epsilon}$ be the standard convolution kernel on $\mathbb{T}^{2}.$ For any real-valued function $f$, we define $f^{(n)}$ by $f^{(n)}:=f\ast\eta_{1/n}$. Since $\Lambda^{-1}\theta^{(n)}_{i}$ can be taken as a test function for all $n\geqslant 1$ and $i=1,2,$ it follows that
$\quad\int_{0}^{t}\left(-\langle\Lambda^{-1/2}\theta_{2},\partial_{t}\Lambda^{-1/2}\theta_{1}^{(n)}\rangle+\langle\Lambda^{\alpha-1/2}\theta_{2},\Lambda^{\alpha-1/2}\theta_{1}^{(n)}\rangle+\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}\theta_{1}^{(n)}\rangle\right)\,\mathrm{d}s$
$=-\langle\Lambda^{-1/2}\theta_{2}(t),\Lambda^{-1/2}\theta_{1}^{(n)}(t)\rangle+\langle\Lambda^{-1/2}\theta_{2}(0),\Lambda^{-1/2}\theta_{1}^{(n)}(0)\rangle+\int_{0}^{t}\langle\Lambda^{\alpha-1/2}\theta_{2},g\rangle\,\mathrm{d}s,$
and
$\quad\int_{0}^{t}\left(-\langle\Lambda^{-1/2}\theta_{1},\partial_{t}\Lambda^{-1/2}\theta_{2}^{(n)}\rangle+\langle\Lambda^{\alpha-1/2}\theta_{1},\Lambda^{\alpha-1/2}\theta_{2}^{(n)}\rangle+\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}\theta_{2}^{(n)}\rangle\right)\,\mathrm{d}s$
$=-\langle\Lambda^{-1/2}\theta_{1}(t),\Lambda^{-1/2}\theta_{2}^{(n)}(t)\rangle+\langle\Lambda^{-1/2}\theta_{1}(0),\Lambda^{-1/2}\theta_{2}^{(n)}(0)\rangle+\int_{0}^{t}\langle\Lambda^{\alpha-1/2}\theta_{1},g\rangle\,\mathrm{d}s.$
It suffices to pass to the limit $n\rightarrow\infty$ to get (A.3). Using Fubini’s theorem and the symmetry of the convolution kernel, we have
$\int_{0}^{t}$ $\langle\Lambda^{-1/2}\theta_{2},\partial_{t}\Lambda^{-1/2}\theta_{1}^{(n)}\rangle\,\mathrm{d}s+\int_{0}^{t}\langle\Lambda^{-1/2}\theta_{1},\partial_{t}\Lambda^{-1/2}\theta_{2}^{(n)}\rangle\,\mathrm{d}s$
$=\langle\Lambda^{-1/2}\theta_{2}(t),\Lambda^{-1/2}\theta_{1}^{(n)}(t)\rangle-\langle\Lambda^{-1/2}\theta_{2}(0),\Lambda^{-1/2}\theta_{1}^{(n)}(0)\rangle.$
Now we focus on the convergence of the nonlinear terms. Since $\theta_{1}\in\mathcal{C}_{0}^{1}$,
$\quad\left|\int_{0}^{t}\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}(\theta_{1}-\theta^{(n)}_{1})\rangle\,\mathrm{d}s\right|$ (A.4)
$\lesssim\int_{0}^{t}\|\theta_{2}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\|\nabla\Lambda^{-1}(\theta_{1}-\theta^{(n)}_{1})\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|u_{\theta_{2}}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2};\mathbb{R}^{2})}\,\mathrm{d}s$
$\lesssim\int_{0}^{t}\|\theta_{1}-\theta^{(n)}_{1}\|_{L^{p}(\mathbb{T}^{2})}\|\theta_{2}\|_{H^{1/p}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$
$\lesssim\int_{0}^{t}\|\theta_{1}-\theta^{(n)}_{1}\|_{L^{p}(\mathbb{T}^{2})}\|\theta_{2}\|_{H^{-1/2}(\mathbb{T}^{2})}^{2/r}\left\|\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2(1-1/r)}\,\mathrm{d}s$
$\lesssim\left(\int_{0}^{t}\|\theta_{1}-\theta^{(n)}_{1}\|_{L^{p}(\mathbb{T}^{2})}^{r}\left\|\theta_{2}\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)^{1/r}\left(\int_{0}^{t}\left\|\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)^{1-1/r}$
$\lesssim\sup_{t\in[0,T]}\left\|\theta_{2}(t)\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2/r}\|\theta_{1}-\theta^{(n)}_{1}\|_{L^{r}\left([0,T];L^{p}(\mathbb{T}^{2})\right)}\left(\int_{0}^{t}\left\|\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)^{1-1/r}\rightarrow 0.$
Here we apply Hölder’s inequality in the second line, Calderón-Zygmund inequality and Sobolev embedding theorem in the third line, Gagliardo-Nirenberg interpolation inequality in the fourth line and Hölder’s inequality in the fifth line. Similarly,
$\quad\bigg{|}\int_{0}^{t}\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}(\theta_{2}-\theta^{(n)}_{2})\rangle\,\mathrm{d}s\bigg{|}\lesssim\int_{0}^{t}\|\theta_{1}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\|u_{\theta_{1}}\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\nabla\Lambda^{-1}(\theta_{2}-\theta^{(n)}_{2})\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\,\mathrm{d}s$
$\lesssim\int_{0}^{t}\|\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}\|\theta_{1}\|_{H^{-1/2}(\mathbb{T}^{2})}^{1/r}\|\theta_{2}-\theta^{(n)}_{2}\|_{H^{-1/2}(\mathbb{T}^{2})}^{1/r}\left\|\theta_{1}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{1-1/r}\left\|\theta_{2}-\theta^{(n)}_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{1-1/r}\,\mathrm{d}s$
$\lesssim\sup_{t\in[0,T]}\left\|\theta_{1}(t)\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{1/r}\sup_{t\in[0,T]}\left\|\theta_{2}(t)-\theta^{(n)}_{2}(t)\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{1/r}\|\theta_{1}\|_{L^{r}\left([0,T];L^{p}(\mathbb{T}^{2})\right)}$
$\quad\cdot\left(\int_{0}^{t}\left\|\theta_{1}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)^{\frac{1-1/r}{2}}\left(\int_{0}^{t}\left\|\theta_{2}-\theta^{(n)}_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)^{\frac{1-1/r}{2}}\rightarrow 0.$
This completes the proof of (A.3).
Since (A.4) implies that $\theta_{1}$ is sufficiently regular to be taken as a test function, for any $t\in[0,T],$
$\frac{1}{2}\|\theta_{1}(t)\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\theta_{1}(s)\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\|\theta_{1}(0)\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\langle\Lambda^{\alpha-1/2}\theta_{1},g\rangle\,\mathrm{d}s.$ (A.5)
According to (A.2), (A.3), (A.5) and (A.1), we obtain
$\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}+\int_{0}^{t}\left\|\theta_{1}-\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ (A.6)
$\leqslant$ $-\int_{0}^{t}\left\|\theta_{1}-\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s+2\int_{0}^{t}\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}\theta_{1}\rangle+\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}\theta_{2}\rangle\,\mathrm{d}s.$
Using the fact that
$\int_{\mathbb{T}^{2}}\theta(u_{\theta_{2}}-u_{\theta_{1}})\cdot\nabla\Lambda^{-1}(\theta_{2}-\theta_{1})\,\mathrm{d}x=0,$
a rearrangement of the terms yields
$\int_{0}^{t}\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}\theta_{1}\rangle+\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}\theta_{2}\rangle\,\mathrm{d}s=\int_{0}^{t}\langle(\theta_{2}-\theta_{1})u_{\theta_{1}},\nabla\Lambda^{-1}(\theta_{2}-\theta_{1})\rangle\,\mathrm{d}s.$ (A.7)
Substituting (A.7) into (A.6), and using the estimate
$\quad\left|\int_{0}^{t}\langle(\theta_{2}-\theta_{1})u_{\theta_{1}}\cdot\nabla\Lambda^{-1}(\theta_{2}-\theta_{1})\rangle\,\mathrm{d}s\right|$ (A.8)
$\leqslant C\int_{0}^{t}\|\theta_{2}-\theta_{1}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\|u_{\theta_{1}}\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\nabla\Lambda^{-1}(\theta_{2}-\theta_{1})\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\,\mathrm{d}s$
$\leqslant C\int_{0}^{t}\|\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}\|\theta_{2}-\theta_{1}\|_{H^{1/p}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$
$\leqslant C\int_{0}^{t}\|\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}\|\theta_{2}-\theta_{1}\|_{H^{-1/2}(\mathbb{T}^{2})}^{2/r}\left\|\theta_{2}-\theta_{1}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2(1-1/r)}\,\mathrm{d}s$
$\leqslant\frac{1}{2}\int_{0}^{t}\left\|\theta_{2}-\theta_{1}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s+C\int_{0}^{t}\|\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}^{r}\left\|\theta_{2}-\theta_{1}\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s,$
we conclude that
$\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}+\int_{0}^{t}\left\|\theta_{1}-\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\lesssim\int_{0}^{t}\|\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}^{r}\left\|\theta_{2}-\theta_{1}\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
It follows from Grönwall’s inequality that $\theta_{1}-\theta_{2}=0.$
@@
In the next example, we will take $\mathcal{C}^{1}:=\mathbb{X}_{\alpha,\alpha/2+1/4}\cap L^{4}\left([0,T];L^{4}(\mathbb{T}^{2})\right)$ and show that $\mathcal{C}^{1}$ is an $\mathcal{I}$-closure of $\mathcal{C}_{0}^{1}.$
@@
Let $\theta\in\mathcal{C}^{1}$ with $\mathcal{I}(\theta)<\infty$. Then there exists a sequence $\{\theta^{(n)}\}_{n\geqslant 1}\subset\mathcal{C}_{0}^{1}$ such that $\theta^{(n)}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\alpha/2+1/4}$ and $\mathcal{I}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}(\theta)$ as $n\rightarrow\infty$.
@@
For every $n\geqslant 1,$ let $\theta^{(n)}:=\theta\ast\eta_{1/n}$. It suffices to prove
(i) $\theta^{(n)}\in\mathcal{C}_{0}^{1}$,
(ii) $\theta^{(n)}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\alpha/2+1/4}$ as $n\rightarrow\infty,$
(iii) $\mathcal{I}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}(\theta)$.
@@
Using Cauchy-Schwartz inequality and Calderón-Zygmund inequality, we have
$\|u_{\theta}\theta\|_{L^{2}\left([0,T];L^{2}(\mathbb{T}^{2};\mathbb{R}^{2})\right)}$ $\lesssim\|u_{\theta}\|_{L^{4}\left([0,T];L^{4}(\mathbb{T}^{2};\mathbb{R}^{2})\right)}\|\theta\|_{L^{4}\left([0,T];L^{4}(\mathbb{T}^{2})\right)}$
$\lesssim\|\theta\|_{L^{4}\left([0,T];L^{4}(\mathbb{T}^{2})\right)}^{2}.$
Since $\alpha>1/2$ and $(u_{\theta}\theta)^{(n)}\rightarrow u_{\theta}\theta$ in $L^{2}\left([0,T];L^{2}(\mathbb{T}^{2};\mathbb{R}^{2})\right)$, we find that
@@
The fact that $\frac{1}{4}+\frac{\alpha}{4}\geqslant\alpha-\frac{1}{2}$ implies that the space $\mathcal{C}^{1}$ is larger than $\mathcal{C}_{0}^{1}$ when taking $r=4$ in the definition of $\mathcal{C}_{0}^{1}$.
@@
Suppose that $20<\alpha\leqslant 1/$ 2 and $\beta=\alpha/2$. Let
$\mathcal{C}_{0}^{2}=\{\theta\in\mathbb{X}_{\alpha,\alpha/2}:\nabla\theta\in L^{r}\left([0,T];L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})\right),\,\frac{1}{p}+\frac{\alpha}{r}=\alpha,\,p>\frac{1}{\alpha},\,r>0\}.$

### v3

@@ 8.2. Potential relationships to open problems for PDEs
Let $\{\nu_{\varepsilon,\delta(\varepsilon)}\}_{\varepsilon>0}$ be a family of ergodic invariant measures of (1.3). Inspired by [BC17] and [CP22], the large deviations for $\{\nu_{\varepsilon,\delta(\varepsilon)}\}_{\varepsilon>0}$ under a scaling regime $(\varepsilon,\delta(\varepsilon))\rightarrow(0,0)$ strongly rely on uniform dynamical large deviations, and its rate function is expected to be governed by the quasi-potential. Based on the discussion in Remark 8.4, one can only expect to observe the rate function for the lower bound governed by the restricted quasi-potential (8.4). This connects to the open problem of uniqueness for the SQG equation (8.5).
As long as one can prove large deviations for $\{\nu_{\varepsilon,\delta(\varepsilon)}\}_{\varepsilon>0}$ with rate function (8.4), then Proposition 8.3 indicates the following relationships. On the one hand, if a solution of (8.5) exists in $\mathcal{A}_{0}$ (which implies uniqueness), then, by applying Proposition 8.3, one can observe that the expected rate function (8.4) equals to the Gaussian rate function $\|\cdot\|^{2}_{H^{\alpha-2\beta}(\mathbb{T}^{2})}$. Conversely, if non-Gaussian large deviations can be proven for $\{\nu_{\varepsilon,\delta(\varepsilon)}\}_{\varepsilon>0}$ in the sense of Definition 8.5, the uniqueness problem of (8.5) could be answered negatively. Therefore, Proposition 8.3 suggests a new perspective for understanding the uniqueness problem of the PDE (8.5) through probabilistic approaches. However, rigorously establishing this relationship requires proving large deviations of the invariant measure, which lies beyond the scope of this paper.
@@
It is worth pointing out that, in [Tot20], the existence of Gaussian stationary solutions for (1.3) with $\delta=0$ is proved using the energy solutions approach. However, this does not contradict the discussion of non-Gaussian large deviations for invariant measures in the sense of Definition 8.5, as the aforementioned non-Gaussian large deviations focus on fluctuations in spaces with higher regularity due to the regularization of the noise and the choice of the scaling regime $(\varepsilon,\delta(\varepsilon))$.
On the other hand, in [HLZZ24], the authors proved there exist infinitely many non-Gaussian ergodic stationary solutions of (1.3) with $\delta=0$ by showing the $L^{\infty}(\Omega\times[0,\infty))$-boundedness of the solutions with respect to some Besov norm. In their work, the parameters $\alpha$ and $\beta$ satisfy $\alpha\in[0,3/4)$ and $\beta\in[0,1/8)$, and the solutions lie in the Besov space $B^{-1/2}_{p,1}$ for any $p\geqslant 2.$ We note that the non-Gaussian large deviations defined above differs from the non-Gaussianity in [HLZZ24]. Definition 8.5 only captures the asymptotic behavior of the measures by analyzing the rate function, demonstrating that it differs from the Gaussian measure $\mathcal{G}\left(0,\varepsilon Q_{\delta(\varepsilon)}/2\right)$ on $H^{\alpha-2\beta}(\mathbb{T}^{2})$. However, it does not exclude the possibility that this family of measures are Gaussian distribution concentrated on a different space. We hope to bridge the discussion between non-Gaussian ergodic stationary solutions and the uniqueness problem of the PDE (8.5) via the non-Gaussian large deviations.
@@ Appendix A A Examples of the Weak-strong Uniqueness Classes
In this section, we provide some examples of the weak-strong uniqueness classes $\mathcal{C}_{0}$ and $\mathcal{R}\subset\overline{\mathcal{C}_{0}}^{\mathcal{I}}$ for different choices of $\alpha$ and $\beta$ satisfying (1.2). These examples cover subcritical, critical, and supercritical cases.
@@
Suppose that $1/2<\alpha<1$ and $\beta=\alpha/2+1/4$. Let
$\mathcal{C}_{0}^{1}:=\{\theta\in\mathbb{X}_{\alpha,\alpha/2+1/4}:\theta\in L^{r}\left([0,T];L^{p}(\mathbb{T}^{2})\right),\,\frac{1}{p}+\frac{\alpha}{r}=\alpha-\frac{1}{2},\,p\geqslant 1,\,r>0\}.$
@@
$\frac{1}{2}\|\theta_{2}(t)\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\|\theta_{2}(s)\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\theta_{2}(0)\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\langle\Lambda^{\alpha-1/2}\theta_{2},g\rangle\,\mathrm{d}s.$ (A.1)
@@
The same conclusion holds in the case where $\beta=\alpha/2$. We refer readers to [CW99_Weak_strong, Theorem 2.2, Remark 2.3] for the proof. ∎
@@
In the next example, we will take $\mathcal{R}^{1}:=\mathbb{X}_{\alpha,\alpha/2+1/4}\cap L^{4}\left([0,T];L^{4}(\mathbb{T}^{2})\right)$ and show that $\mathcal{R}^{1}$ is contained in the $\mathcal{I}$-closure of $\mathcal{C}_{0}^{1}.$
@@
Let $\theta\in\mathcal{R}^{1}$ with $\mathcal{I}(\theta)<\infty$. Then there exists a sequence $\{\theta^{(n)}\}_{n\geqslant 1}\subset\mathcal{C}_{0}^{1}$ such that $\theta^{(n)}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\alpha/2+1/4}$ and $\mathcal{I}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}(\theta)$ as $n\rightarrow\infty$.
@@
Let $\{\eta_{\epsilon}\}_{\epsilon\in(0,1)}$ be the standard convolution kernel on $\mathbb{T}^{2}.$ For any real-valued function $f$ and every $n\geqslant 1$ , we define $f^{(n)}$ by $f^{(n)}:=f\ast\eta_{1/n}$. Now for $\theta^{(n)}:=\theta\ast\eta_{1/n}$, it suffices to prove
(i) $\theta^{(n)}\in\mathcal{C}_{0}^{1}$, (ii) $\theta^{(n)}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\alpha/2+1/4}$ as $n\rightarrow\infty,$ (iii) $\mathcal{I}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}(\theta)$.
@@
Using Cauchy-Schwarz inequality, Calderón-Zygmund inequality, and the fact that $\theta\in\mathcal{R}^{1},$ we have $u_{\theta}\theta\in L^{2}\left([0,T];L^{2}(\mathbb{T}^{2};\mathbb{R}^{2})\right)$. Since $\alpha>1/2$ and $(u_{\theta}\theta)^{(n)}\rightarrow u_{\theta}\theta$ in $L^{2}\left([0,T];L^{2}(\mathbb{T}^{2};\mathbb{R}^{2})\right)$, we find that
@@
The fact that $\frac{1}{4}+\frac{\alpha}{4}\geqslant\alpha-\frac{1}{2}$ implies that the space $\mathcal{R}^{1}$ is larger than $\mathcal{C}_{0}^{1}$ when taking $r=4$ in the definition of $\mathcal{C}_{0}^{1}$.
@@
Suppose that $0<\alpha\leqslant 1/2$ and $\beta=\alpha/2$. Let
$\mathcal{C}_{0}^{2}:=\{\theta\in\mathbb{X}_{\alpha,\alpha/2}:\nabla\theta\in L^{r}\left([0,T];L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})\right),\,\frac{1}{p}+\frac{\alpha}{r}=\alpha,\,p>\frac{1}{\alpha},\,r>0\}.$